Phantomdex:

Quadrilateral RSTU, diagonals SU and RT intersect at point V. RSTU is a parallelogram. If m∠TUV = 78° and m∠TVU = 54°, explain how you can find the measure of ∠SRV. Show all steps of your work, and refer to any properties of triangles, parallelograms, or triangle congruency theorems as necessary to justify your response.

Phantomdex:

AZ:

Do you know what the sum of angles in a triangle is?

Phantomdex:

Um well, I have no clue what's going on in this question. I don't know what to do on it at all nor how to solve it.

AZ:

In a triangle, what do all the angles add up to?

Phantomdex:

Shouldn't they add up to a degree of 180?

AZ:

Yes! Can you find the third angle in the triangle now?

1 attachment
Phantomdex:

Is it 45?

AZ:

How did you get that?

AZ:

All three angles should add up to 180. x + 54 + 78 = 180 x = ??

Phantomdex:

Yes, so shouldn't x = 45 because 54 and 78 give us 132 and if you add 45 that equals up to 180

Phantomdex:

If I'm doing it correctly that is

AZ:

No 132 + 45 = 177

AZ:

To find the value faster, you can just do 180 - 132 = ?

Phantomdex:

Oh, right I completely forgot about that method

Phantomdex:

Is it 48 then?

AZ:

Now that you found the angle of VTU, let's find the value of angle SRV Do you know the alternate interior angle theorem?

Phantomdex:

Isn't that when two parallel lines are cut by transversal?

AZ:

Essentially, the alternate interior angle theorem says that angle VTU and SRV are congruent because the two parallel sides are cut by a transversal line. Thus, the alternate interior angles are equal to each other

AZ:

You got it!

Phantomdex:

Oh, nice!

Phantomdex:

So, what about SRV?

AZ:

It is equal to angle VTU which we calculated as 48 degrees

Phantomdex:

Yes, but I need to be able to explain how I've found the measure of SRV. Also needing to show my work. Unless we've done that already.

AZ:

We've done it all in two steps

Phantomdex:

Oh, then I apologize for being difficult.

AZ:

1) You used the sum of all angles in a triangle is 180 2) You used the alternate interior angles theorem

AZ:

@phantomdex wrote:
Oh, then I apologize for being difficult.
No worries! Geometry proofs can be a hassle

Phantomdex:

Indeed